OpenStudy (anonymous):

Find four consecutive intergers such that the product of the 1st and 3rd is 25 greater than of -13 and the fourth

OpenStudy (mathteacher1729):

Weinburgerd, what part of this problem is giving you difficulty? I can help in a bunch of ways, but It helps if I know what's giving you trouble. :)

OpenStudy (anonymous):

Ok Its been a while since i have taken math and I was asked to help. This is what we came up with so far.

OpenStudy (anonymous):

n n-1 n-2 n-3

OpenStudy (anonymous):

n(n-2)=25-13(n-3) n^2 -2n= 25-13n +39 n^2 - 2n= 64 -13n n^2= 64-11n how do we find what n equals?

OpenStudy (anonymous):

I set the four consecutive integers to n, n+1,n+2 & n+3 n(n+2)=25-13(n+3) n^2 -2n= 25-13n -39 n^2 - 2n= -13n-14 n^2= -15n-14 n^2+15n+14=0 use the quadratic equation to solve n=-14 or -1

OpenStudy (anonymous):

Oops cut and paste and forgot to change the sign n^2 + 2n= 25-13n -39 n^2 + 2n= -13n-14 n^2= -15n-14 n^2+15n+14=0 use the quadratic equation to solve n=-14 or -1